Michael is 6 years older than Monique. In 3 years the sum of their ages will be 54. How old is Michael now?
Answer by josgarithmetic(39617)

Michael age x,
Monique age y.
"Today": x=y+6
"In three years...": (x+3)+(y+3)=54

Try substituting x from Today into the threemoreyears equation.
(y+6+3)+(y+3)=54
y+9+y+3-54=0
2y+12-54=0
y+6-27=0
y-21=0
y=21. x=21+6=27